OpenStudy (anonymous):

-5(3-a)=-3(2-a) Solve for the variable. Steps please

OpenStudy (anonymous):

Start by distributing the numbers outside the parenthesis to the terms inside the parenthesis.

OpenStudy (anonymous):

Let me know what you get when you do that.

OpenStudy (anonymous):

-13-5a=-6-3a ??

OpenStudy (anonymous):

Okay so this is when you have to multiply everything out. It's called the Distributive Property, and it means that no matter what order you multiply anything, your result will be the same. So you multiply -5 by both 3 and -a (think of 3 and -a as two separate values) and for the other side of the equation, multiply -3 by both 2 and -a. This gives you:\[-5(3-a)=-3(2-a)\]\[-15+5a=-6+3a\] Where you seem to be having trouble is with the negatives -- remember that two negatives multiplied together equal a positive! So -5*-a will give you a positive 5a, and the same thing goes for -3 and -a. So now you're only half done. You need to actually find the value of a, and that looks like: \[-15+15+5a=-6+15+36\]\[5a=9+3a\]\[5a-3a=9+3a-3a\]\[2a=9\]\[a=4.5\] And that's your answer!